A remark on 3-D Navier-Stokes equations

Luigi Carlo Berselli, Dipartimento di matematica Università di Pisa Via Buonarroti 2 56127 Pisa Italy

                  berselli@dm.unipi.it

We find a sufficient condition that ensures the regularity of weak solutions to the Navier-Stokes equations. Our conditions involve the pressure and can be seen at the light of some heuristic ideas that link the behavior of velocity and pressure.